# Token-level Accept or Reject: A Micro Alignment Approach for Large Language Models

### v2

@@ 4.3 Ablation Studies
Reward distribution SafeRLHF BeaverTails HarmfulQA
@@
Mistral-7B-v0.1 $\alpha_{r}$:$\alpha_{c}$=1:0 +29.65% -26.13% +2.01% +37.29% -22.57% +7.29% +21.93% -18.44% +1.43%
$\alpha_{r}$:$\alpha_{c}$=0:1 -55.78% +63.82% +4.02% -34.29% +51.57% +8.57% -49.39% +53.89% +2.46%
$\alpha_{r}$:$\alpha_{c}$=1:1 -37.49% +49.25% +6.03% -19.71% +48.71% +14.43% -40.78% +51.84% +5.53%
$\alpha_{r}$:$\alpha_{c}$=2:1 -16.58% +28.69% +11.06% -2.71% +37.21% +17.43% -22.75% +43.44% +10.45%
Mistral-7B-v0.2 $\alpha_{r}$:$\alpha_{c}$=1:0 +24.62% -20.60% +2.01% +31.71% -21.14% +5.29% +27.05% -22.95% +2.05%
$\alpha_{r}$:$\alpha_{c}$=0:1 -91.96% +40.70% -1.01% -30.00% +33.71% +1.86% -52.87% +52.46% -0.20%
$\alpha_{r}$:$\alpha_{c}$=1:1 -22.61% +35.68% +6.53% -15.43% +35.14% +9.71% -41.19% +50.41% +4.51%
$\alpha_{r}$:$\alpha_{c}$=2:1 -20.6% +30.15% +4.52% +9.43% +19.29% +14.29% -31.56% +41.6% +4.92%
Mistral-7B-v0.3 $\alpha_{r}$:$\alpha_{c}$=1:0 +29.65% -9.05% +10.55% +37.57% -11.29% +13.00% +11.07% -9.84% +0.61%
$\alpha_{r}$:$\alpha_{c}$=0:1 -39.70% +65.83% +13.07% -37.71% +45.71% +4.00% -56.97% +60.25% +1.64%
$\alpha_{r}$:$\alpha_{c}$=1:1 -31.66% +54.27% +11.56% -17.14% +43.14% +13.00% -50.41% +69.06% +9.22%
$\alpha_{r}$:$\alpha_{c}$=2:1 -17.39% +50.75% +17.09% -2.29% +34.71% +16.00% -42.62% +64.75% +11.07%
Table 5: Impact of reward signal distribution on model performance in terms of helpful, harmless, and preference rate. $\alpha_{r}$:$\alpha_{c}$ represents the ratio between reward model (beaver-7b-v1.0-reward) and cost model (beaver-7b-v1.0-cost) signals in the reward of $r(x,y)$. Bold numbers indicate the best performance under each metric (helpful/harmless/preference) for each model version.
To analyze the impact of different reward signal distributions, we conduct experiments with various ratios between reward and cost model signals ($\alpha_{r}$:$\alpha_{c}$). As shown in Table 5, we observe distinct trade-offs between helpfulness and harmlessness across different configurations. When solely using the reward model ($\alpha_{r}$:$\alpha_{c}$=1:0), the model shows improved helpfulness (up to +37.57%) but decreased harmlessness. Conversely, using only the cost model ($\alpha_{r}$:$\alpha_{c}$=0:1) leads to significant gains in harmlessness (up to +65.83%) but often at the expense of helpfulness. A balanced ratio of $\alpha_{r}$:$\alpha_{c}$=2:1 achieves the best overall performance across datasets, with preference rate improvements reaching +17.09% on SafeRLHF, +16.00% on BeaverTails, and +11.07% on HarmfulQA using Mistral-7B-v0.3. This suggests that while both reward signals are important, slightly emphasizing the reward model over the cost model leads to optimal balance between helpfulness and safety. Detailed ablation studies on the impact of candidate token set size and the sensitivity to KL divergence coefficients are presented in Appendix B.2 and Appendix B.3, respectively.

### v3

@@ 4.3 Ablation Studies
Reward
distribution
SafeRLHF BeaverTails HarmfulQA
@@
Mistral-7B-v0.1 $\alpha_{r}$$\alpha_{c}$=1:0 +29.65% -26.13% +2.01% +37.29% -22.57% +7.29% +21.93% -18.44% +1.43%
$\alpha_{r}$$\alpha_{c}$=0:1 -55.78% +63.82% +4.02% -34.29% +51.57% +8.57% -49.39% +53.89% +2.46%
$\alpha_{r}$$\alpha_{c}$=1:1 -37.49% +49.25% +6.03% -19.71% +48.71% +14.43% -40.78% +51.84% +5.53%
$\alpha_{r}$$\alpha_{c}$=2:1 -16.58% +28.69% +11.06% -2.71% +37.21% +17.43% -22.75% +43.44% +10.45%
Mistral-7B-v0.2 $\alpha_{r}$$\alpha_{c}$=1:0 +24.62% -20.60% +2.01% +31.71% -21.14% +5.29% +27.05% -22.95% +2.05%
$\alpha_{r}$$\alpha_{c}$=0:1 -91.96% +40.70% -1.01% -30.00% +33.71% +1.86% -52.87% +52.46% -0.20%
$\alpha_{r}$$\alpha_{c}$=1:1 -22.61% +35.68% +6.53% -15.43% +35.14% +9.71% -41.19% +50.41% +4.51%
$\alpha_{r}$$\alpha_{c}$=2:1 -20.6% +30.15% +4.52% +9.43% +19.29% +14.29% -31.56% +41.6% +4.92%
Mistral-7B-v0.3 $\alpha_{r}$$\alpha_{c}$=1:0 +29.65% -9.05% +10.55% +37.57% -11.29% +13.00% +11.07% -9.84% +0.61%
$\alpha_{r}$$\alpha_{c}$=0:1 -39.70% +65.83% +13.07% -37.71% +45.71% +4.00% -56.97% +60.25% +1.64%
$\alpha_{r}$$\alpha_{c}$=1:1 -31.66% +54.27% +11.56% -17.14% +43.14% +13.00% -50.41% +69.06% +9.22%
$\alpha_{r}$$\alpha_{c}$=2:1 -17.39% +50.75% +17.09% -2.29% +34.71% +16.00% -42.62% +64.75% +11.07%
Table 5: Impact of reward signal distribution on model performance in terms of helpful, harmless, and preference rate. $\alpha_{r}$$\alpha_{c}$ represents the ratio between reward model (beaver-7b-v1.0-reward) and cost model (beaver-7b-v1.0-cost) signals in the reward of $r(x,y)$. Bold numbers indicate the best performance under each metric (helpful/harmless/preference) for each model version.
To analyze the impact of different reward signal distributions, we conduct experiments with various ratios between reward and cost model signals ($\alpha_{r}$$\alpha_{c}$). As shown in Table 5, we observe distinct trade-offs between helpfulness and harmlessness across different configurations. When solely using the reward model ($\alpha_{r}$$\alpha_{c}$=1:0), the model shows improved helpfulness (up to +37.57%) but decreased harmlessness. Conversely, using only the cost model ($\alpha_{r}$$\alpha_{c}$=0:1) leads to significant gains in harmlessness (up to +65.83%) but often at the expense of helpfulness. A balanced ratio of $\alpha_{r}$$\alpha_{c}$=2:1 achieves the best overall performance across datasets, with preference rate improvements reaching +17.09% on SafeRLHF, +16.00% on BeaverTails, and +11.07% on HarmfulQA using Mistral-7B-v0.3. This suggests that while both reward signals are important, slightly emphasizing the reward model over the cost model leads to optimal balance between helpfulness and safety. Detailed ablation studies on the impact of candidate token set size and the sensitivity to KL divergence coefficients are presented in Appendix B.2 and Appendix B.3, respectively.